OpenStudy (anonymous):

4+3s-9s=22 a.s=-3 b.s=-1.5 c.s=1.5 d.s=4/3

OpenStudy (anonymous):

First,subtract 4 on both sides

OpenStudy (anonymous):

\[4-4+3s-9s=22-4\]\[-6s=18\]

MsBrains (ms-brains):

Simplify is easy, all you have to do is work your way down to the shortest answer.

OpenStudy (anonymous):

Then,divide both sides by -6

OpenStudy (anonymous):

\[\frac{ -6s }{ -6 }=\frac{ 18 }{ -6 }\]

OpenStudy (anonymous):

\[s=?\]solve it @angelpooh01